OpenStudy (anonymous):

the flower shop sells it's own brand of fertilizer it mixes one brand containing 7% of nitrogen with another that contains 8% of nitrogen. the resulting mixture has 7.7% nitrogen. how much of the 8% fertilizer must be used to make 200lbs of the 7.7% mixture?

OpenStudy (anonymous):

please help me...

OpenStudy (amistre64):

a(.07) + b(.08) = 200(.077)

OpenStudy (amistre64):

since a+b = 200; we say that a=200-b is a good value to plug in

OpenStudy (amistre64):

(200-b)(.07) + b(.08) = 200(.077) 200(.07) -.07b +.08b = 200(.077)

OpenStudy (amistre64):

.01 b = 200(.077) - 200(.07) b = 200(.077 -.07) ------------ .01

OpenStudy (amistre64):

when you know the value for b, use it to find a by: 200 - b = a

OpenStudy (anonymous):

thankss so much for helping me because i"m trying to finsh school. and it's really helpful that your giving me the steps.:) your a angel from god.

OpenStudy (amistre64):

well, thnx. your welcome

OpenStudy (amistre64):

you dont have to find the value for a; since b is the 8% anyways; and you might wanna dbl chk this but I get 140

OpenStudy (anonymous):

it was right thaank you:) the book store sells their own brand of coffee. they mix 120lbs of the mixture to seel each day. the mixture of columbain coffee sells for $5/lb and the brazilian coffee $8/lb. the mixture sells for a total of $7.00 per lb. hany many lbs are needed of the brailian do they need to make 120lbs?

OpenStudy (anonymous):

can someone help me please

OpenStudy (amistre64):

c + b = 120; c = 120 -b c(5) + b(8) = 120(7) (120-b)(5) + b(8) = 120(7) ; they want brazilians, so solve for b

OpenStudy (anonymous):

thanks. when you give me the equations i can figure them out;) sliced turkey costs $3.20/lb. roast beef costs twice as much, the ground beef costs half as turkey. find the price of 1.75lb turkey

OpenStudy (amistre64):

t = 3.20 rb = 2(3.20) = 6.40 gb = (3.20)/2 = 1.60 the price for 1.75lbs turkey is the question? (turkey price per pound)(lbs used) 3.20(1.75) if i read the question right; kinda confused why the put all that other stuff in there.

OpenStudy (anonymous):

idk either but the answer is right;)

OpenStudy (amistre64):

post new answers to the left of your screen to get a better chance of people seeing them :)